# Polar Mounds on Strangeon Stars: the Neutrino Emission from Ultraluminous X-ray Pulsars

### v1

@@ 5 Summary and discussion
We find that at relatively low mass accretion rates ($10^{20}\,\rm g\,s^{-1}$), photons can escape freely from the accretion column, and the total luminosity is dominated by photon emission, while neutrino luminosity remains negligible. At higher mass accretion rates ($10^{21}\,\rm g\,s^{-1}$), photon trapping becomes significant, and energy is transported to regions cooled by neutrinos, resulting in a total luminosity that can exceed the photon luminosity. For extremely high accretion rates, most of the accretion energy is converted into neutrinos, leading to a saturation of the photon luminosity around $10^{41}\,\rm erg\,s^{-1}$. Furthermore, the luminosity associated with heat transport along the SS surface is estimated to be $L_{\rm heat}\sim 1.2\times 10^{36}\,\rm erg\,s^{-1}$, independent of the mass accretion rate.
Based on our model, we calculated the expected neutrino fluxes from several known ULXPs. For most extragalactic sources, the predicted neutrino flux at Earth is extremely low due to large distances, with the notable exception of Swift J0243.6++6124, the closest ULXP, where the neutrino flux could reach $F_{\nu}\sim 2.3\times 10^{-2}\,\rm MeV\,cm^{-2}\,s^{-1}$. We compared these fluxes with the expected neutrino background, including contributions from old supernovae, atmospheric neutrinos, and geoneutrinos. Our results indicate that even for the closest source, detection with current neutrino observatories would be challenging.
This study highlights the importance of considering the detailed structure of the thermal mound and EOS of SSs in modeling high-luminosity accretion flows. While photon emission dominates at low accretion rates, neutrino emission can become the primary cooling channel at extreme accretion rates, potentially affecting the interpretation of observed ULXP luminosities. Future work could include multidimensional modeling of the accretion column, detailed neutrino spectra, and the effects of neutrino oscillations, which may provide more accurate predictions for potential detection by next-generation neutrino telescopes.

### v2

@@ 5 Summary and discussion
We find that at relatively low mass accretion rates ($10^{20}\,\rm g\,s^{-1}$), photons can escape freely from the accretion column, and the total luminosity is dominated by photon emission, while neutrino luminosity remains negligible. At higher mass accretion rates ($10^{21}\,\rm g\,s^{-1}$), photon trapping becomes significant, and energy is transported to regions cooled by neutrinos, resulting in a total luminosity that can exceed the photon luminosity. To compare with the NS model, we also compute the neutrino luminosity from the accretion column using Eq. (1) of Asthana et al. (2023). At relatively low accretion rates ($10^{20}\,\rm g\,s^{-1}$), the SS model predicts moderately higher neutrino fluxes for Swift J0243.6++6124, RX J0209.6-7427, and SMC X-3, exceeding the results of Asthana et al. (2023) by factors of $\sim 2-5$. In contrast, for ULXPs, the SS model yields neutrino fluxes that are comparable to or lower than those of the NS model, with typical reductions of $\lesssim 2$ and up to a factor of $\sim 5$ for extreme sources such as NGC 5907 ULX-1. Furthermore, the luminosity associated with heat transport along the SS surface is estimated to be $L_{\rm heat}\sim 1.2\times 10^{36}\,\rm erg\,s^{-1}$, independent of the mass accretion rate.
Within the framework of the SS model, our results indicate that the predicted neutrino flux at Earth is extremely low due to large distances for most extragalactic sources, with the notable exception of Swift J0243.6++6124, the closest ULXP, where the neutrino flux could reach $F_{\nu}\sim 2.3\times 10^{-2}\,\rm MeV\,cm^{-2}\,s^{-1}$. However, even under optimistic parameter choices, the predicted fluxes remain below the diffuse MeV neutrino background. Related discussions of neutrino emission from accreting neutron stars can also be found in the literature, for example in Asthana et al. (2023).
Concerning prospects for point-source detection, the quantitative results shown in Fig. 5 indicate that the expected neutrino fluxes from Swift J0243.6++6124, RX J0209.6-7427, NGC 5907 ULX-1, and M51 ULX-8 typically lie one to two orders of magnitude below the diffuse background level in the relevant energy range. Therefore, even for the closest and brightest systems currently known, the signal would be strongly background-dominated. A rough implication of our results is that only sources at distances of order $\lesssim 1\,\rm kpc$, and with comparably high luminosities, might approach the threshold of potential detectability, although such a detection remains highly challenging with present instruments.
This study highlights the importance of considering the detailed structure of the thermal mound and EOS of SSs in modeling high-luminosity accretion flows. While photon emission dominates at low accretion rates, neutrino emission can become the primary cooling channel at extreme accretion rates, potentially affecting the interpretation of observed ULXP luminosities.
Acknowledgements
@@
Data Availability
The data underlying this paper will be shared on reasonable request to the corresponding authors.